• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name gilman

    Log entry time 16:02:07 on May 9,2002

    Entry number 3016

    keyword=shift summary


    8:00 waiting for fast feedback to come back on

    8:15 hear no FFB until this afternoon, move to optics target

    8:25 ask for beam to be centered

    8:28 run 1382, optics target, 70 uA, 2x2 raster on

    8:35 switch to thin C target, have to wait for magnets afterward

    8:53 run 1383, thin C target, 70 uA, raster on, beam center ~0.1,0.0

    9:09 run 1384, thin C target, 70 uA, raster off, beam center ~0.1,0.0
    with FFB off, beam seems to wander by 0.1 at times.

    before changing angles,
    front back
    left 36.5 deg, -11.2 mm 36.25 deg, +8.1 mm
    right 65.5 deg, -5.7 mm 65.5 deg, +29.4 mm

    9:24 start angle change, afterwards
    front back
    left 40.25 deg, +3.8 mm 40.25 deg, -39.5 mm
    40.5 deg, -39.0 mm 40.5 deg, +30.3 mm

    9:38 run 1385, thin C target,... beam off

    10:08 beam off ~35 min, troubleshooting ffb, intermittent fault found,
    maybe another 20 minutes.

    10:17 MCC calls and says ready to send us beam

    10:33 beam tuned, run 1386, thin C, raster off, centered at 0.,0.
    10:45 run 1387, thin C, raster on, centered at 0.0, 0.0
    11:04 run 1388, LH2, production data
    11:22 run 1389, LH2, production data
    11:45 run 1390, LH2, production data
    12:08 run 1391, LH2, production data
    12:22 ask for beam to be taken away to set up for access
    for spectrometer survey

    14:15 run 1392, cosmics run (to check detectors)

    14:30 beam back!

    14:30 run 1396, LH2, production data
    14:55 run 1397, LH2, production data
    switch to dummy target
    15:36 run 1398, dummy data
    lower current to 30 uA! Ooopps.
    15:55 run 1399, dummy data at 30 uA.
    - end of shift -